# SDEs with supercritical distributional drifts

### v1

@@ 5.1. Existence of weak solutions
where $b_{n}=b*\phi_{n}$. Recall the definition of ${\mathscr{I}}_{d}$ in (3.11) and the parameter set $\Theta$ in (3.3). We first show the following uniform Krylov estimate.
@@
$\sup_{n\in{\mathbb{N}}}{\mathbf{E}}\left(\int^{\tau_{1}}_{\tau_{0}}f(s,X^{n}_{s}){\mathord{{\rm d}}}s\Big{|}{\mathscr{F}}_{\tau_{0}}\right)\lesssim_{C}\|f\|_{{\mathbb{L}}^{q}_{t}{\mathbf{H}}^{\alpha}_{p}}.$
@@
$\sup_{n}\Big{(}\|u_{n}\|_{{\mathbb{L}}^{\infty}_{t}}+\|u_{n}\|_{{\mathscr{V}}_{t}}\Big{)}\lesssim_{C}\|f\|_{{\mathbb{L}}^{q}_{t}{\mathbf{H}}^{\alpha}_{p}}.$ (5.2)
@@
Let ${\mathscr{T}}_{T}$ be the set of all stopping times bounded by $T$. By Aldous’ criterion of tightness, it suffices to prove that for some increasing function $\ell:(0,1)\to(0,\infty)$ with $\lim_{\delta\to 0}\ell(\delta)=0$,
$\sup_{\tau,\eta\in{\mathscr{T}}_{T},\eta\leqslant\tau\leqslant\eta+\delta}\sup_{n}{\mathbf{E}}|X_{\tau}^{n}-X_{\eta}^{n}|\leqslant\ell(\delta),\ \ \delta\in(0,1),$
where $\ell(\delta)$ will be determined below. By the strong Markov property, we only need to show
$\sup_{t\in[0,T]}\sup_{x_{0}\in{\mathbb{R}}^{d}}\sup_{\tau\leqslant\delta}\sup_{n}{\mathbf{E}}|X_{\tau}^{t,n}(x_{0})-x_{0}|\leqslant\ell(\delta),$
@@
Define
$h_{\delta}(x):=\sqrt{\ell(\delta)^{2}/9+|x-x_{0}|^{2}}.$
@@
${\mathbf{E}}|X^{t,n}_{\tau}(x_{0})-x_{0}|\leqslant{\mathbf{E}}h_{\delta}(X^{t,n}_{\tau}(x_{0}))=\ell(\delta)/3+{\mathbf{E}}\left(\int^{\tau}_{0}(\Delta+b_{n}(s)\cdot\nabla)h_{\delta}(X^{t,n}_{s}(x_{0})){\mathord{{\rm d}}}s\right).$
Note that for some $C$ independent of $\delta$ and $t,x_{0}$,
$|\nabla h_{\delta}(x)|=\frac{|x-x_{0}|}{\sqrt{\ell(\delta)^{2}/9+|x-x_{0}|^{2}}}\leqslant 1,\ \ |\nabla^{2}h_{\delta}|\leqslant 2\ell(\delta)^{-1}.$ (5.4)
@@
${\mathbf{E}}\left(\int^{\tau}_{0}\Delta h_{\delta}(X^{t,n}_{s}(x_{0})){\mathord{{\rm d}}}s\right)\leqslant\delta\|\Delta h_{\delta}\|_{\infty}\leqslant 2d\delta/\ell(\delta).$
On the other hand, since $\frac{d}{p_{b}}+\frac{2}{q_{b}}<2+\alpha$, without loss of generality we assume $q_{b}<\infty$. Otherwise, we may choose $q_{b}^{\prime}<\infty$ so that $\frac{d}{p_{b}}+\frac{2}{q^{\prime}_{b}}<2+\alpha$. By Lemma 5.1 and (iv) of Lemma 2.4, we have
${\mathbf{E}}\left(\int^{\delta\wedge\tau}_{0}(b^{t}_{n}(s)\cdot\nabla h_{\delta})(X^{t,n}_{s}(x_{0})){\mathord{{\rm d}}}s\right)$ $\leqslant C_{1}\|b^{t}_{n}\cdot\nabla h_{\delta}\|_{{\mathbb{L}}^{q_{b}}_{\delta}{\mathbf{H}}^{\alpha_{b}}_{p_{b}}}\leqslant C_{2}\|b^{t}_{n}\|_{{\mathbb{L}}^{q_{b}}_{\delta}{\mathbf{H}}^{\alpha_{b}}_{p_{b}}}\|\nabla h_{\delta}\|_{C^{1}_{b}},$
where the constant $C_{2}>0$ only depends on $\Theta$. By (5.4), we further have
${\mathbf{E}}\left(\int^{\delta\wedge\tau}_{0}(b^{t}_{n}(s)\cdot\nabla h_{\delta})(X^{t,n}_{s}(x_{0})){\mathord{{\rm d}}}s\right)\leqslant C_{3}\|b^{t}\|_{{\mathbb{L}}^{q_{b}}_{\delta}{\mathbf{H}}^{\alpha_{b}}_{p_{b}}}/\ell(\delta)\leqslant(C_{3}\|b^{t}\|_{{\mathbb{L}}^{q_{b}}_{\delta}{\mathbf{H}}^{\alpha_{b}}_{p_{b}}}\vee(6\delta))/\ell(\delta).$
Finally if we take
$\ell(\delta):=\sup_{t\in[0,T]}\sqrt{3[C_{3}\|b^{t}\|_{{\mathbb{L}}^{q_{b}}_{\delta}{\mathbf{H}}^{\alpha_{b}}_{p_{b}}}\vee(6d\delta)]},$
then combining the above calculations, we obtain
${\mathbf{E}}|X^{n}_{\tau}(x_{0})-x_{0}|\leqslant\ell(\delta)/3+2d\delta/\ell(\delta)+\ell(\delta)/3\leqslant\ell(\delta).$
The proof is complete since $\lim_{\delta\to 0}\sup_{t\in[0,T]}\int^{(t+\delta)\wedge T}_{t}|g(s)|{\mathord{{\rm d}}}s=0$ for $g\in L^{1}([0,T])$. ∎
@@
$\|\rho_{n}\|_{{\mathscr{V}}_{T}}\leqslant(\kappa^{b}_{2}+1)\mathrm{e}^{\kappa^{b}_{2}}\|\rho_{0}\|_{2},$ (5.6)
and
@@
First of all, since $b_{n}\in{\mathbb{L}}^{2}_{T}C^{\infty}_{b}$, by the maximum principle, we have
$\rho_{n}\in{\mathbb{L}}^{\infty}_{T}L^{\infty}.$
@@
$\frac{1}{2}\partial_{t}\|\rho_{n}\varphi\|_{2}^{2}+\|\nabla(\rho_{n}\varphi)\|_{2}^{2}=\langle b_{n}\rho_{n}\varphi_{n},\nabla(\rho_{n}\varphi)\rangle+\langle g^{\varphi}_{n},\rho_{n}\varphi\rangle.$
@@
$\langle b_{n}\rho_{n}\varphi_{n},\nabla(\rho_{n}\varphi)\rangle=-\frac{1}{2}\langle\mathord{{\rm div}}b_{n},(\rho_{n}\varphi)^{2}\rangle\leqslant\frac{1}{2}\|\mathord{{\rm div}}b_{n}\|_{\infty}\|\rho_{n}\varphi\|^{2}_{2}\leqslant\frac{1}{2}\|\mathord{{\rm div}}b\|_{\infty}\|\rho_{n}\varphi\|^{2}_{2}$
@@
$\|\rho_{n}\varphi\|_{{\mathscr{V}}_{T}}\leqslant(\kappa^{b}_{2}+1)\mathrm{e}^{\kappa^{b}_{2}}\left(\|\rho_{0}\varphi\|_{2}+\|\langle b_{n},\varphi\rho^{2}_{n}\nabla\varphi\rangle\|_{L^{1}_{T}}+\|\langle\rho^{2}_{n},|\nabla\varphi|^{2}\rangle\|_{L^{1}_{T}}\right).$ (5.9)
@@
$\|\rho_{n}\|_{{\mathscr{V}}_{T}}\leqslant(\kappa^{b}_{2}+1)\mathrm{e}^{\kappa^{b}_{2}}\|\rho_{0}\|_{2}.$
@@
$\|\rho_{n}\bar{\chi}_{R}\|_{{\mathscr{V}}_{T}}\leqslant(\kappa^{b}_{2}+1)\mathrm{e}^{\kappa^{b}_{2}}\left(\|\rho_{0}\bar{\chi}_{R}\|_{2}+\|\langle b_{n},\bar{\chi}_{R}\rho^{2}_{n}\nabla\bar{\chi}_{R}\rangle\|_{L^{1}_{T}}+\|\langle\rho^{2}_{n},|\nabla\bar{\chi}_{R}|^{2}\rangle\|_{L^{1}_{T}}\right).$
@@
${\mathbf{E}}\left|\int^{t}_{0}f(s,X^{n}_{s}){\mathord{{\rm d}}}s\right|^{2}\lesssim_{C}\|f\|^{2}_{{\mathbb{L}}^{q}_{t}{\mathbf{H}}^{\alpha}_{p}}\|\rho_{0}\|_{2}.$ (5.10)
@@
One can not show the following stronger estimate
${\mathbf{E}}\left(\sup_{t\in[0,T]}\left|\int^{t}_{0}f(s,X^{n}_{s}){\mathord{{\rm d}}}s\right|^{2}\right)\lesssim_{C}\|f\|^{2}_{{\mathbb{L}}^{q}_{T}{\mathbf{H}}^{\alpha}_{p}}\|\rho_{0}\|_{2}$
@@
${\mathbf{E}}\left|\int^{t}_{0}f(s,X_{s}){\mathord{{\rm d}}}s\right|^{2}\lesssim_{C}\|f\|^{2}_{{\mathbb{L}}^{q}_{t}{\mathbf{H}}^{\alpha}_{p}}\|\rho_{0}\|_{2},$ (5.12)
where $C=C(\Theta,\alpha,p,q)>0$. Moreover, for each $t\in(0,T]$, $X_{t}$ admits a density $\rho_{t}\in L^{2}$ with $\|\rho\|_{{\mathscr{V}}_{T}}\leqslant(\kappa^{b}_{2}+1)\mathrm{e}^{\kappa^{b}_{2}}\|\rho_{0}\|_{2}$ and in the distributional sense
@@
By Lemma 5.3 and Prokhorov’s criterion, there is a subsequence $n_{k}$ and a probability measure ${\mathbb{P}}\in{\mathcal{P}}({\mathbb{C}}_{T})$ so that ${\mathbb{P}}_{n_{k}}$ weakly converges to ${\mathbb{P}}$ as $k\to\infty$. Without loss of generality, we assume that ${\mathbb{P}}_{n}$ weakly converges to ${\mathbb{P}}$ as $n\to\infty$. By Skorokhod’s representation theorem, if necessary, changing the probability space, we may assume that there is a continuous process $(X_{t})_{t\in[0,T]}$ such that
@@
Estimate (5.12) follows by taking limits for (5.10). To show that $X$ is a weak solution of SDE (1.1), it suffices to show that
$\lim_{n\to\infty}{\mathbf{E}}\left|\int_{0}^{t}b_{n}(r,X^{n}_{r}){\mathord{{\rm d}}}r-\int_{0}^{t}b_{n}(r,X_{r}){\mathord{{\rm d}}}r\right|^{2}=0.$
Noticing that for fixed $m\in{\mathbb{N}}$, by the dominated convergence theorem,
@@
once we have shown
@@
then it follows that
$\lim_{n\to\infty}\int_{0}^{t}b_{n}(r,X_{r}){\mathord{{\rm d}}}r=:A^{b}_{t}\mbox{ in $L^{2}$}.$
@@ 5.2. Generalized martingale problems
Let $(\alpha,p,q)\in{\mathscr{I}}_{d}$ (see (3.11) for a definition). For any $f\in{\mathbb{L}}^{q}_{T}{\mathbf{H}}^{\alpha}_{p}$, by Theorem 3.6, there exists a weak solution to the following backward PDE
@@
Note that the weak solution may be not unique if we do not assume $b\in{\mathbb{L}}^{\infty}_{T}{\mathcal{B}}+{\mathbb{L}}^{2}_{T}{\mathbb{L}}$. However, if we consider the following regularized PDE
@@
$\|{\mathcal{S}}_{f}^{n}\|_{{\mathbb{L}}^{\infty}_{T}\cap{\mathscr{V}}_{T}}\leqslant C\|f\|_{{\mathbb{L}}^{q}_{T}{\mathbf{H}}^{\alpha}_{p}},$
where $C$ is independent of $n$. In other words, $f\mapsto{\mathcal{S}}_{f}^{n}$ is a bounded linear operator from ${\mathbb{L}}^{q}_{T}{\mathbf{H}}^{\alpha}_{p}$ to ${\mathbb{L}}^{\infty}_{T}\cap{\mathscr{V}}_{T}$. By Banach-Steinhaus theorem, there is a subsequence $n_{k}$ and a bounded linear operator ${\mathcal{S}}:{\mathbb{L}}^{q}_{T}{\mathbf{H}}^{\alpha}_{p}\to{\mathbb{L}}^{\infty}_{T}\cap{\mathscr{V}}_{T}$ so that for each $f\in{\mathbb{L}}^{q}_{T}{\mathbf{H}}^{\alpha}_{p}$ and $t\in[0,T]$, $\varphi\in L^{1}({\mathbb{R}}^{d})$ (see the proof of Theorem 3.6),
@@
In the following we fix such an operator ${\mathcal{S}}$ and construct a unique generalized martingale solution associated with ${\mathcal{S}}$. Of course, its definition depends on the choice of subsequence $n_{k}$ as well as the mollifiers $\phi_{n_{k}}$. We emphasize that when $b\in{\mathbb{L}}^{\infty}_{T}{\mathcal{B}}+{\mathbb{L}}^{2}_{T}{\mathbb{L}}$, ${\mathcal{S}}_{f}$ is nothing, but the unique solution of (5.16), which is independent of the choice of $n_{k}$ and the mollifiers.
@@
Let $\mu\in{\mathcal{P}}({\mathbb{R}}^{d})$. We call a probability measure ${\mathbb{P}}\in{\mathcal{P}}({\mathbb{C}}_{T})$ a generalized martingale solution of SDE (1.1) starting from $\mu$ and associated with the operator ${\mathcal{S}}$, if ${\mathbb{P}}\circ({w}_{0})^{-1}=\mu$ and for any $f\in C^{\infty}_{c}([0,T]\times{\mathbb{R}}^{d})$,
$M_{t}^{f}:={\mathcal{S}}_{f}(t,{w}_{t})-{\mathcal{S}}_{f}(0,{w}_{0})-\int_{0}^{t}f(r,{w}_{r}){\mathord{{\rm d}}}r,\ \ w_{\cdot}\in{\mathbb{C}}_{T},$
is an almost surely martingale under ${\mathbb{P}}$ with respect to the natural filtration ${\mathscr{B}}_{s}$ in the sense that there is a Lebesgue null set ${\mathscr{N}}\subset(0,T)$ such that for all $0\leqslant s<t\leqslant T$ with $s,t\notin{\mathscr{N}}$,
@@
Since, for each $t\in[0,T]$, ${\mathcal{S}}_{f}(t,\cdot)\in L^{\infty}$, in order to make sense of $M_{t}^{f}$, we have assumed that the marginal distribution ${\mathbb{P}}\circ w^{-1}_{t}$ has a density. However, it’s essential to note that although $t\mapsto M^{f}_{t}$ is defined pointwise, we lack information about the path properties of $t\mapsto M^{f}_{t}$. For instance, we have no knowledge of its cádlág property since we only have ${\mathcal{S}}_{f}\in{\mathbb{L}}^{\infty}_{T}$.
@@
(Uniqueness) Let ${\mathbb{P}}_{1},{\mathbb{P}}_{2}$ be two generalized martingale solutions with initial distribution $\mu_{0}({\mathord{{\rm d}}}x)=\rho_{0}(x){\mathord{{\rm d}}}x$ in the sense of Definition 5.8. By the definition, for any $f\in C_{c}([0,T]\times{\mathbb{R}}^{d})$, there is a Lebesgue full measure set ${\mathscr{A}}_{f}\subset[0,T]$ containing 00 and $T$ such that for any $s,t\in{\mathscr{A}}_{f}$ and bounded ${\mathscr{B}}_{s}$-measurable $G_{s}$,
@@
${\mathbb{E}}^{{\mathbb{P}}_{1}}\left(\int_{0}^{T}f(t,{w}_{t}){\mathord{{\rm d}}}t\right)=\langle{\mathcal{S}}_{f}(0),\rho_{0}\rangle={\mathbb{E}}^{{\mathbb{P}}_{2}}\left(\int_{0}^{T}f(t,{w}_{t}){\mathord{{\rm d}}}t\right).$

### v3

@@ 5.1. Existence of weak solutions
where $b_{n}=b*\phi_{n}$. Recall the definition of ${\mathbf{I}}_{d}$ in (3.10) and the parameter set $\Theta$ in (3.2). We first show the following uniform Krylov estimate.
@@
$\sup_{n\in{\mathbb{N}}}{\mathbf{E}}\left(\int^{\tau_{1}}_{\tau_{0}}f(s,X^{n}_{s}){\mathord{{\rm d}}}s\Big{|}{\mathscr{F}}_{\tau_{0}}\right)\leqslant C\|f\|_{{\mathbb{L}}^{q}_{t}{\mathbf{H}}^{\alpha}_{p}}.$
@@
$\sup_{n}\Big{(}\|u_{n}\|_{{\mathbb{L}}^{\infty}_{t}}+\|u_{n}\|_{{\mathscr{V}}_{t}}\Big{)}\leqslant C\|f\|_{{\mathbb{L}}^{q}_{t}{\mathbf{H}}^{\alpha}_{p}}.$ (5.2)
@@
Without loss of generality, we assume that $q_{b}<\infty$. In the case where $q_{b}=\infty$ and $d/p_{b}<2+\alpha_{b}$, one can choose $\tilde{q}_{b}<\infty$ such that $2/\tilde{q}_{b}+d/p_{b}<2+\alpha_{b}$ and $\kappa_{b}=\|b\|_{{\mathbb{L}}^{\tilde{q}_{b}}{\mathbf{H}}^{\alpha_{b}}_{p_{b}}}<\infty$.
Let ${\mathscr{T}}_{T}$ be the set of all stopping times bounded by $T$. By Aldous’ criterion of tightness, it suffices to prove that
$\lim_{\delta\downarrow 0}\sup_{\tau,\eta\in{\mathscr{T}}_{T},\eta\leqslant\tau\leqslant\eta+\delta}\sup_{n}{\mathbf{E}}|X_{\tau}^{n}-X_{\eta}^{n}|=0.$
By the strong Markov property, we only need to show
$\lim_{\delta\downarrow 0}\sup_{t\in[0,T]}\sup_{x_{0}\in{\mathbb{R}}^{d}}\sup_{\tau\leqslant\delta}\sup_{n}{\mathbf{E}}|X_{\tau}^{t,n}(x_{0})-x_{0}|=0,$
@@
For $\varepsilon\in(0,1)$, define
$h_{\varepsilon}(x):=\sqrt{\varepsilon^{2}+|x-x_{0}|^{2}}.$
@@
${\mathbf{E}}|X^{t,n}_{\tau}(x_{0})-x_{0}|\leqslant{\mathbf{E}}h_{\varepsilon}(X^{t,n}_{\tau}(x_{0}))=\varepsilon+{\mathbf{E}}\left(\int^{\tau}_{0}(\Delta+b_{n}^{t}(s)\cdot\nabla)h_{\varepsilon}(X^{t,n}_{s}(x_{0})){\mathord{{\rm d}}}s\right).$
Note that for some $C$ independent of $\varepsilon$ and $t,x_{0}$,
$|\nabla h_{\varepsilon}(x)|=\frac{|x-x_{0}|}{\sqrt{\varepsilon^{2}+|x-x_{0}|^{2}}}\leqslant 1,\ \ |\nabla^{2}h_{\varepsilon}|\leqslant 2d^{2}/\varepsilon.$ (5.4)
@@
${\mathbf{E}}\left(\int^{\tau}_{0}\Delta h_{\varepsilon}(X^{t,n}_{s}(x_{0})){\mathord{{\rm d}}}s\right)\leqslant\delta\|\Delta h_{\varepsilon}\|_{\infty}\leqslant 2d\delta/\varepsilon.$
Moreover, by Lemma 5.1, (iv) of Lemma 2.4 and (5.4), we have
${\mathbf{E}}\left(\int^{\delta\wedge\tau}_{0}(b^{t}_{n}(s)\cdot\nabla h_{\varepsilon})(X^{t,n}_{s}(x_{0})){\mathord{{\rm d}}}s\right)\leqslant C_{1}\|b^{t}_{n}\cdot\nabla h_{\varepsilon}\|_{{\mathbb{L}}^{q_{b}}_{\delta}{\mathbf{H}}^{\alpha_{b}}_{p_{b}}}$
$\qquad\leqslant C_{2}\|b^{t}_{n}\|_{{\mathbb{L}}^{q_{b}}_{\delta}{\mathbf{H}}^{\alpha_{b}}_{p_{b}}}\|\nabla h_{\varepsilon}\|_{C^{1}_{b}}\leqslant C_{3}\|b^{t}\|_{{\mathbb{L}}^{q_{b}}_{\delta}{\mathbf{H}}^{\alpha_{b}}_{p_{b}}}/\varepsilon.$
Combining the above calculations, we obtain
${\mathbf{E}}|X^{t,n}_{\tau}(x_{0})-x_{0}|\leqslant\varepsilon+2d\delta/\varepsilon+C_{3}\|b^{t}\|_{{\mathbb{L}}^{q_{b}}_{\delta}{\mathbf{H}}^{\alpha_{b}}_{p_{b}}}/\varepsilon.$
Since $\lim_{\delta\to 0}\sup_{t\in[0,T]}\int^{(t+\delta)\wedge T}_{t}\|b(s)\|^{q_{b}}_{{\mathbf{H}}^{\alpha_{b}}_{p_{b}}}{\mathord{{\rm d}}}s=0,$ firstly letting $\delta\to 0$ and then $\varepsilon\to 0$, we complete the proof. ∎
@@
$\|\rho_{n}\|_{{\mathscr{V}}_{T}}\leqslant(\tilde{\kappa}_{b}+1)\mathrm{e}^{\tilde{\kappa}_{b}}\|\rho_{0}\|_{2},$ (5.6)
where ${\mathscr{V}}_{T}$ is defined in (3.11), and
@@
First of all, since $\mathord{{\rm div}}b_{n}\in{\mathbb{L}}^{2}_{T}C^{\infty}_{b}$ and $\rho_{0}\in L^{2}$, it is standard to derive that
$\rho_{n}\in{\mathbb{L}}^{\infty}_{T}L^{2}.$
@@
$\frac{1}{2}\partial_{t}\|\rho_{n}\varphi\|_{2}^{2}+\|\nabla(\rho_{n}\varphi)\|_{2}^{2}=\langle b_{n}\rho_{n}\varphi,\nabla(\rho_{n}\varphi)\rangle+\langle g^{\varphi}_{n},\rho_{n}\varphi\rangle.$
@@
$\langle b_{n}\rho_{n}\varphi,\nabla(\rho_{n}\varphi)\rangle=-\frac{1}{2}\langle\mathord{{\rm div}}b_{n},(\rho_{n}\varphi)^{2}\rangle\leqslant\frac{1}{2}\|\mathord{{\rm div}}b_{n}\|_{\infty}\|\rho_{n}\varphi\|^{2}_{2}\leqslant\frac{1}{2}\|\mathord{{\rm div}}b\|_{\infty}\|\rho_{n}\varphi\|^{2}_{2}$
@@
$\|\rho_{n}\varphi\|_{{\mathscr{V}}_{T}}\leqslant(\tilde{\kappa}_{b}+1)\mathrm{e}^{\tilde{\kappa}_{b}}\left(\|\rho_{0}\varphi\|_{2}+\|\langle b_{n},\varphi\rho^{2}_{n}\nabla\varphi\rangle\|_{L^{1}_{T}}+\|\langle\rho^{2}_{n},|\nabla\varphi|^{2}\rangle\|_{L^{1}_{T}}\right).$ (5.9)
@@
$\|\rho_{n}\|_{{\mathscr{V}}_{T}}\leqslant(\tilde{\kappa}_{b}+1)\mathrm{e}^{\tilde{\kappa}_{b}}\|\rho_{0}\|_{2}.$
@@
$\|\rho_{n}\bar{\chi}_{R}\|_{{\mathscr{V}}_{T}}\leqslant(\tilde{\kappa}_{b}+1)\mathrm{e}^{\tilde{\kappa}_{b}}\left(\|\rho_{0}\bar{\chi}_{R}\|_{2}+\|\langle b_{n},\bar{\chi}_{R}\rho^{2}_{n}\nabla\bar{\chi}_{R}\rangle\|_{L^{1}_{T}}+\|\langle\rho^{2}_{n},|\nabla\bar{\chi}_{R}|^{2}\rangle\|_{L^{1}_{T}}\right).$
@@
${\mathbf{E}}\left|\int^{t}_{0}f(s,X^{n}_{s}){\mathord{{\rm d}}}s\right|^{2}\leqslant C\|f\|^{2}_{{\mathbb{L}}^{q}_{t}{\mathbf{H}}^{\alpha}_{p}}\|\rho_{0}\|_{2}.$ (5.10)
@@
We are not able to show the following stronger estimate
${\mathbf{E}}\left(\sup_{t\in[0,T]}\left|\int^{t}_{0}f(s,X^{n}_{s}){\mathord{{\rm d}}}s\right|^{2}\right)\leqslant C\|f\|^{2}_{{\mathbb{L}}^{q}_{T}{\mathbf{H}}^{\alpha}_{p}}\|\rho_{0}\|_{2}$
@@
${\mathbf{E}}\left|\int^{t}_{0}f(s,X_{s}){\mathord{{\rm d}}}s\right|^{2}\leqslant C\|f\|^{2}_{{\mathbb{L}}^{q}_{t}{\mathbf{H}}^{\alpha}_{p}}\|\rho_{0}\|_{2},$ (5.12)
where $C=C(\Theta,\alpha,p,q)>0$. Moreover, for each $t\in(0,T]$, $X_{t}$ admits a density $\rho_{t}\in L^{2}$ with
$\|\rho\|_{{\mathscr{V}}_{T}}\leqslant(\tilde{\kappa}_{b}+1)\mathrm{e}^{\tilde{\kappa}_{b}}\|\rho_{0}\|_{2}$ (5.13)
and in the distributional sense
@@
By Lemma 5.3 and Prokhorov’s criterion, there exists a subsequence $n_{k}$ and a probability measure ${\mathbb{P}}\in{\mathcal{P}}({\mathbb{C}}_{T})$ such that ${\mathbb{P}}_{n_{k}}$ weakly converges to ${\mathbb{P}}$ as $k\to\infty$. Without loss of generality, we assume that ${\mathbb{P}}_{n}$ weakly converges to ${\mathbb{P}}$ as $n\to\infty$. By Skorokhod’s representation theorem, if necessary, changing the probability space, we may assume that there exists a continuous process $(X_{t})_{t\in[0,T]}$ such that
@@
Estimate (5.12) follows by taking limits in (5.10). To show that $X$ is a weak solution of SDE (1.1), it suffices to show that
$\lim_{n\to\infty}{\mathbf{E}}\left|\int_{0}^{t}b_{n}(r,X^{n}_{r}){\mathord{{\rm d}}}r-\int_{0}^{t}b_{n}(r,X_{r}){\mathord{{\rm d}}}r\right|^{2}=0.$ (5.15)
Notice that for fixed $m\in{\mathbb{N}}$, by the dominated convergence theorem,
@@
and once we have shown
@@
then (5.15) follows. Moreover, we also have
$\lim_{n\to\infty}\int_{0}^{t}b_{n}(r,X_{r}){\mathord{{\rm d}}}r=:A^{b}_{t}\quad\text{in $L^{2}$}.$
@@ 5.2. Generalized martingale problems
Let $(\alpha,p,q)\in{\mathbf{I}}_{d}$ (see (3.10) for a definition). For any $f\in{\mathbb{L}}^{q}_{T}{\mathbf{H}}^{\alpha}_{p}$, by Theorem 3.6, there exists a weak solution to the following backward PDE
@@
Note that the weak solution may be not unique if we do not assume $b\in{\mathbb{L}}^{\infty}_{T}{\mathcal{B}}+{\mathbb{L}}^{2}_{T}L^{2}$. However, if we consider the following regularized PDE
@@
$\|{\mathcal{S}}_{f}^{n}\|_{{\mathbb{L}}^{\infty}_{T}\cap{\mathscr{V}}_{T}}\leqslant C\|f\|_{{\mathbb{L}}^{q}_{T}{\mathbf{H}}^{\alpha}_{p}},$ (5.20)
where $C$ is independent of $n$. In other words, $f\mapsto{\mathcal{S}}_{f}^{n}$ is a bounded linear operator from ${\mathbb{L}}^{q}_{T}{\mathbf{H}}^{\alpha}_{p}$ to ${\mathbb{L}}^{\infty}_{T}\cap{\mathscr{V}}_{T}$.
Now suppose $p,q<\infty$. Let $\{f_{m}\}_{m=1}^{\infty}\subset{\mathbb{L}}^{q}_{T}{\mathbf{H}}^{\alpha}_{p}$ be a countable dense subset. By the Cantor-Hilbert selection and from the proof of Theorem 3.6, there is a subsequence $\{n_{k}\}$ such that for each $f=f_{m}$, there is a solution ${\mathcal{S}}_{f}=:u$ to PDE (5.18) such that for any $t\in[0,T]$ and $\varphi\in L^{1}({\mathbb{R}}^{d})$,
@@
Moreover, by (5.20) and (5.21), we have
$\|{\mathcal{S}}_{f_{m}}\|_{{\mathbb{L}}^{\infty}_{T}\cap{\mathscr{V}}_{T}}\leqslant C\|f_{m}\|_{{\mathbb{L}}^{q}_{T}{\mathbf{H}}^{\alpha}_{p}},\quad m\in{\mathbb{N}}.$
Thus ${\mathcal{S}}_{f}$ can be extended as a bounded linear operator from ${\mathbb{L}}^{q}_{T}{\mathbf{H}}^{\alpha}_{p}$ to ${\mathbb{L}}^{\infty}_{T}\cap{\mathscr{V}}_{T}$ so that (5.21) and (5.22) holds for all $f\in{\mathbb{L}}^{q}_{T}{\mathbf{H}}^{\alpha}_{p}$.
In the following we fix such an operator ${\mathcal{S}}$ and construct a unique generalized martingale solution associated with ${\mathcal{S}}$. Of course, its definition depends on the choice of subsequence $n_{k}$ as well as the mollifiers $\phi_{n_{k}}$. We emphasize that when $b\in{\mathbb{L}}^{\infty}_{T}{\mathcal{B}}+{\mathbb{L}}^{2}_{T}L^{2}$, ${\mathcal{S}}_{f}$ is nothing, but the unique solution of (5.18), which is independent of the choice of $n_{k}$ and the mollifiers.
@@
Let $\mu\in{\mathcal{P}}({\mathbb{R}}^{d})$. We call a probability measure ${\mathbb{P}}\in{\mathcal{P}}({\mathbb{C}}_{T})$ a generalized martingale solution of SDE (1.1) starting from $\mu$ and associated with the operator ${\mathcal{S}}$, if ${\mathbb{P}}\circ({w}_{0})^{-1}=\mu$, and for each $t\in(0,T)$, ${\mathbb{P}}\circ w_{t}^{-1}$ has a density with respect to the Lebesgue measure, and for any $f\in C^{\infty}_{c}([0,T]\times{\mathbb{R}}^{d})$,
$M_{t}^{f}:={\mathcal{S}}_{f}(t,{w}_{t})-{\mathcal{S}}_{f}(0,{w}_{0})-\int_{0}^{t}f(r,{w}_{r}){\mathord{{\rm d}}}r,\quad w_{\cdot}\in{\mathbb{C}}_{T},$
is an almost surely martingale under ${\mathbb{P}}$ with respect to the natural filtration ${\mathscr{B}}_{s}$, in the sense that there is a Lebesgue null set ${\mathscr{N}}\subset(0,T)$ such that for all $0\leqslant s<t\leqslant T$ with $s,t\notin{\mathscr{N}}$,
@@
Although $t\mapsto M^{f}_{t}$ is defined pointwise, we lack information about the path properties of $t\mapsto M^{f}_{t}$. For instance, we have no knowledge of its cádlág property since we only have ${\mathcal{S}}_{f}\in{\mathbb{L}}^{\infty}_{T}$.
@@
In particular, by (5.6), for each $t\in(0,T)$, ${\mathbb{P}}\circ(w_{t})^{-1}$ has a density w.r.t. the Lebesgue measure.
@@
(Uniqueness) Let ${\mathbb{P}}_{1},{\mathbb{P}}_{2}$ be two generalized martingale solutions with initial distribution $\mu_{0}({\mathord{{\rm d}}}x)=\rho_{0}(x){\mathord{{\rm d}}}x$ in the sense of Definition 5.8. By the definition, for any $f\in C_{c}([0,T]\times{\mathbb{R}}^{d})$, there is a Lebesgue full measure set ${\mathscr{A}}_{f}\subset[0,T]$ containing 0 and $T$ such that for any $s,t\in{\mathscr{A}}_{f}$ and bounded ${\mathscr{B}}_{s}$-measurable $G_{s}$,
@@
${\mathbb{E}}^{{\mathbb{P}}_{1}}\left(\int_{0}^{T}f(t,{w}_{t}){\mathord{{\rm d}}}t\right)=-\langle{\mathcal{S}}_{f}(0),\rho_{0}\rangle={\mathbb{E}}^{{\mathbb{P}}_{2}}\left(\int_{0}^{T}f(t,{w}_{t}){\mathord{{\rm d}}}t\right).$